OpenStudy (anonymous):

find the x and y intercepts for each following questions 1. 5x-3y-15=0 2. x+2y+3=0 3. -3x+5y-15=0 4. y=6-6x

OpenStudy (anonymous):

X-intercept:when y=0 y-intercept:when x=0 so for the first one x intercept: 5x-3(0)-15=0 5x=15 x=3 y intercept: 5(0)-3y-15=0 -3y=15 y=-5